[Brahms-l] Run status

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Wed May 04 2005 - 19:13:58 EDT
During the todays access a number of fixes and changes was made to the exp setup.

Dana:
1) put tofII onto MRS platform.
2) Moved MRS to 90 degrees
3) Replaced and checked the MRST0 disc.
4) Moved the finger counters (CC response check)
Ramiro+Bob:
   looked at C4/tfw2 HV pods. Identified and fixed problems (we think)
BoB:
fixed 3 mux for DC readiout that was blown arouind Apr 29.
FV:
Included spin scalers into normal start/stop of runs.
Please refer to 
http://server.c-ad.bnl.gov/esfd/RMEM/rhic_planning.htm
for RHIC sand other experiment status.

we are presenttly taking An measurements for pions at 2.3 deg.

Flemming

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed May 4 19:14:48 2005

This archive was generated by hypermail 2.1.8 : Wed May 04 2005 - 19:15:00 EDT